OpenStudy (anonymous):

(64/121) TO THE -½ POWER

OpenStudy (anonymous):

\[(\frac{ 64 }{ 121 })^{-1/2}\]

OpenStudy (anonymous):

You can't use a calculator?

OpenStudy (kkutie7):

^(1/2) is the same as square root

OpenStudy (kkutie7):

^to a negative number means you move it to the denominator. in this case 1/here

zepdrix (zepdrix):

\[\Large\rm \left(\frac{64}{121}\right)^{-1/2}=\left(\frac{121}{64}\right)^{1/2}\]Yah, the negative tells you to flip the fraction :) Numerator becomes denominator and denominator becomes numerator. Then apply what kutie said,\[\Large\rm =\sqrt{\frac{121}{64}}=\frac{\sqrt{121}}{\sqrt{64}}=?\]

OpenStudy (anonymous):

11/8… Thank you!

zepdrix (zepdrix):

Yay good job \c:/